Big Bend , CA

      Site ID:     bbd
      Latitude:     39.309000 N
      Longitude:     -120.518000 W
      Elevation:     1754 meters
      Active:     Y
      Site Table:     Site Table
      Data:     Realtime Display

Instrument Name     Project Name     Install Date     Data Types
Datalogger CR1000     Hydrometeorology Testbed - West     2013-04-15          view
Net Radiation REBS Q7.1     Hydrometeorology Testbed - West     2007-10-22          view
Pressure Sensor - PTB101B     Hydrometeorology Testbed - West     2007-10-22          view
Snow Depth - Judd     Hydrometeorology Testbed - West     2007-10-22          view
Soil Moisture 10cm     Hydrometeorology Testbed - West     2007-10-22          view
Soil Moisture 15cm     Hydrometeorology Testbed - West     2007-10-22          view
Soil Temperature 10cm     Hydrometeorology Testbed - West     2007-10-22          view
Soil Temperature 15cm     Hydrometeorology Testbed - West     2007-10-22          view
Temp/RH HMP45C     Hydrometeorology Testbed - West     2007-10-22          view
Tipping Bucket Heated 385     Hydrometeorology Testbed - West     2007-10-22          view
Wind Speed and Direction 5103     Hydrometeorology Testbed - West     2007-10-22          view

Instrument Name     Project Name     Install Date     Removal Date     Data Types
Datalogger CR23X     Hydrometeorology Testbed - West     2007-10-22     2013-04-15          view
Solar Radiation CM3     Hydrometeorology Testbed - West     2007-10-22     2013-04-15          view
Hotplate Precipitation TPS-3100     Hydrometeorology Testbed - West     2011-11-03     2012-05-08          view
Disdrometer Parsivel Optical     Hydrometeorology Testbed - West     2011-11-03     2012-05-08          view
Disdrometer Parsivel Optical     Hydrometeorology Testbed - West     2012-11-16     2013-04-15          view
Hotplate Precipitation TPS-3100     Hydrometeorology Testbed - West     2012-11-16     2013-04-15          view
Hotplate Precipitation TPS-3100     Hydrometeorology Testbed - West     2013-11-05     2014-04-08          view
Disdrometer Parsivel Optical     Hydrometeorology Testbed - West     2013-11-05     2014-04-08          view
Hotplate Precipitation TPS-3100     Hydrometeorology Testbed - West     2014-12-05     2015-06-01          view
Disdrometer Parsivel Optical     Hydrometeorology Testbed - West     2014-12-05     2015-06-01          view

  OutageID     SiteID     Failure Type     Start Date      End Date      Description     Resolution Description  
724 bbd Communication - Cellular 2020-06-10 13:02:00 Unable to connect to cellular modem
698 bbd Communication - Cellular 2020-03-07 13:00:00 2020-03-13 18:00:00 Can't connect to cellular modem; suspected serial-server crash Datalogger OS updated and cellular modem firmware rolled back to version v.2.008
689 bbd Communication - Cellular 2020-02-06 15:00:00 2020-02-10 21:30:00 Can't connect to cellular modem Modem power source switched from CS/IO port to direct battery connection
637 bbd Hardware - Instrument 2018-06-27 13:18:00 2019-07-19 02:52:00 Datalogger battery failure; 9-volt reading prior to outage The CGFI power switch had tripped at the power box. Reset and installed new battery in the data logger. 7-18-2019
520 bbd Communication - Cellular 2017-03-08 12:00:00 2017-03-08 16:00:00 Datalogger unreachable. Communications restored.
528 bbd Hardware - Instrument 2017-02-24 00:00:00 2017-06-20 18:00:00 Heated tipping bucket precipitation gauge malfunction. Tipping bucket replaced.
468 bbd Hardware - Instrument 2015-10-05 12:00:00 2015-10-20 00:00:00 The datalogger will not run a program. Tried sending a program via logger net remotely and keypad but does not run. Datalogger should be replaced. Datalogger replaced.
448 bbd Hardware - Instrument 2015-04-23 05:00:00 2015-04-27 16:00:00 Unable to talk to datalogger via PC. Could be damaged to lighting in the area. RAD modem was damaged. Replaced.
438 bbd Hardware - Instrument 2015-02-01 00:00:00 2015-03-10 00:00:00 The 10cm and 15cm probes are reporting invalid data. Disabled for display. soil probes fixed.
376 bbd Hardware - Instrument 2014-04-07 00:00:00 2014-05-01 00:00:00 Net Radiation dome broke. Removed from display until fixed. dome replaced.
374 bbd Hardware - Instrument 2014-03-30 00:00:00 2014-05-01 00:00:00 PC needs to replaced. Powered off. Powered down for season. Will replace next season.
298 bbd Hardware - PC 2013-01-13 19:00:00 2013-04-16 18:00:00 Datalogger service stops and hangs intermittently after running for several hours. Investigating causes. Update. 3/4/13. New Service installed but still failing. Computer was scheduled to be replaced but new computer was damaged during shipping. Datalogger changed to CR10x.
285 bbd Hardware - Instrument 2012-12-14 16:00:00 2013-01-07 23:00:00 Snow depth sensor failed. Snow depth sensor replaced.
276 bbd Power Failure 2012-12-02 15:00:00 2012-12-02 19:00:00 Modem unresponsive. Could be a power failure, modem issue, phone line issue, or comms computer issue. Comms restored. Power outage likely.
269 bbd Communication - Phone 2012-10-15 16:00:00 2012-11-30 21:00:00 Modem not answering. Could be power or comms modem issue. Update 10/18/12. Phone line taken down due to construction in the area. Will replace with cellular modem on next visit. Data collecting locally. Replaced modem with Cellular router.
160 bbd Communication - Phone 2011-11-11 00:00:00 2011-11-20 00:00:00 Crossed phone line. Local technician working on the problem. Local contact fixed crossed phone line.
149 bbd Communication - Phone 2011-10-18 22:00:00 2011-11-10 00:00:00 Modem does not output tone. Possibly cause is phone network performing upgrade. NIST phone engineers troubleshooting the problem. According to NIST telecom, the Long Distance carrier made some adjustments to the trunks. Issue should be resolved.
145 bbd Hardware - Instrument 2011-09-26 15:00:00 2011-11-01 00:00:00 Net Radiometer reporting false readings. Removed from display. Will replace on next site visit. Approx late Oct early Nov. Replaced Net Radiometer
92 bbd Communication - Phone 2011-05-08 22:00:00 2011-05-09 13:00:00 Hung modem in laboratory modem pool. Modem hardware reset. Data backfilled.
79 bbd Communication - Phone 2011-03-27 12:00:00 2011-03-30 00:00:00 Modem unresponsive. Could be a power failure, modem issue, phone line issue, or comms computer issue. Modem connection restored. Snow/Ice build up likely cause for outage.
69 bbd Hardware - Instrument 2011-03-19 14:00:00 2011-04-14 22:00:00 Snow-depth sensor is buried in the snow. Last reliable reading was 103 inches measured on 3-19-11. Snow depth reporting realistic snow depths. Snow pack likely melted lower than snow depth sensor.